Mendy was working out on the treadmill. She worked out for 15 minutes. Part of the time she was walking at 3.3 miles per hour. Her heart rate during that time was 108 beats per minute. THe rest of the time she was slowly jogging at 4.2 miles per hour. Her heart rate during that time was 129 beats per minute. During the 15 minutes she covered .89 mile.
How much time did she spend walking and how much time did she spend jogging?
What was her average heart rate during the 15 minutes?

Answer by Paul(988)

Let x be the time of walking
Let y be the time of jogging
EQUATION:
x+y=15
y=15-x (subsitution)
EQUATION 2:
3.3x+4.2y=89
SUbsitute for y and solve
FOR THE AVERAGE HEART RATE:
(129+108)/2=118.5
Paul.